# arXiv:2602.24140v1[physics.ins-det] 27 Feb 2026

### v1

@@ Abstract
This paper presents a radiation-hardened current-mode delta-sigma ADC fabricated in a standard 130 nm CMOS technology and qualified for total ionizing doses up to 100 Mrad. The converter is designed for beam loss monitoring applications in high-energy physics, where it must handle input currents spanning nine decades, from 1 mA down to 1 pA, while providing a fast 10 µs response time for machine protection. To meet these conflicting requirements, the architecture exploits the inherent trade-off between resolution and acquisition time: a first-order modulator sampled at 20 MHz delivers 11-bit effective resolution within the critical 10 µs window for the mA current range. Extended integration times of up to 100 s enable the sub-picoampere resolution required for beam alignment and background monitoring and provides an operational dynamic range exceeding 200 dB. The chip integrates two independent channels, consumes 25 mW from a 1.2 V supply, and includes radiation-hardening techniques such as triple-redundant digital logic and SEU-tolerant comparator banks. Post-irradiation measurements up to 100 Mrad show no performance degradation, and the uncalibrated integral nonlinearity remains within [+0.2%, –0.3%] of full scale over the 1 mA to 5 µA range. The converter’s flexibility and radiation tolerance make it suitable not only for the HL-LHC beam loss monitoring upgrade but also for other precision current measurement applications in harsh environments.
@@ I Introduction
RADIATION-TOLERANT current digitizers primarily find their use in harsh environments where radiation monitoring is required, like in high energy physics experiments and nuclear power plants. For example, at CERN the radiation levels are carefully monitored around the Large Hadron Collider in order to protect the machine components and to accurately position the absorber jaws of the mechanical beam collimation system. In both cases the sensing element is a gas filled ionization chamber [1, 2, 3] that produces a fixed charge rate when subject to constant radiation levels; the corresponding electrical signal ranges from 1 mA to few picoamperes (pA) that represent respectively the maximum tolerable radiation energy, which absorbed within the response time of the beam dumping system, and the typical background radiation levels. Various current sensing interfaces exist either built with discrete components [4] or as integrated circuits [5]; in any case their radiation tolerance limit imposes a lower bound for the length of the cable connecting the interface with the sensor. Long cables limit the detection dynamic range due to interference, noise charge associated with the parasitic capacitance and leakage current. The radiation levels around the LHC are maximum at the collimators preceding the interaction points (IP3 and IP7) [6] and are expected to grow after the High Luminosity upgrade foreseen by 2025. The radiation around the collider originates from the losses of the beams circulating inside the machine, since the relevant quantity is the dose rate absorbed within one turn rather than its instantaneous value, the application band of interest spans from dc up to 12 kHz. An integrated radiation tolerant frontend has therefore been developed to comply with the stringent requirements of radiation monitoring in harsh environments. To this end a standard CMOS 130 nm technology was selected after being successfully characterized to tolerate a total ionizing dose (TID) of 200 Mrad [7, 8, 9] when adopting specific radiation hardness measures to prevent latch up, drain/source leakage and to tolerate SEU and SET in the digital parts. The remainder of this manuscript is organized as follows. Section II presents the system requirements and design drivers derived from the Beam Loss Monitoring application and the HL-LHC upgrade constraints. Section III describes the architecture of the proposed delta-sigma current digitizer, discussing its resolution limit and stability considerations while also providing the foundational concepts for readers less familiar with this architecture. Section IV details the circuit-level implementation, including the integrator, quantizer, adaptive clock control and DAC. Section V reports the experimental results, including noise performance, linearity, and dynamic range as a function of acquisition time, for samples irradiated with X-rays up to 100 Mrad. Finally, Section VI concludes the paper.
@@ II System Requirements and Design Drivers
The Beam Loss Monitoring (BLM) system [10] is a critical protection element of the LHC accelerator complex, designed to prevent quenches of superconducting magnets and equipment damage caused by excessive radiation following beam misalignment. The system continuously measures radiation levels along the accelerator using approximately 4000 ionization chambers installed throughout the 27 km of the circular collider. When ionizing radiation passes through the gas-filled chambers, the resulting charge is extracted by a high-voltage bias of 1.5 kV and converted into a current signal proportional to the instantaneous beam loss. In the original LHC design these signals are transmitted via coaxial cables, up to 800 m in length, to about 750 acquisition modules for processing and analysis.
To ensure compatibility with the higher luminosity and correspondingly increased radiation levels expected from the HL-LHC upgrade, a new generation of front-end electronics must meet stringent performance, reliability, and environmental constraints. The enhanced radiation tolerance enables the readout electronics to be installed closer to the sensors, thereby improving signal integrity. However, depending on the radiation levels and associated aging effects, the electronics must also remain compatible with cable lengths of up to 50 m to preserve installation flexibility, although with relaxed requirements for the non-critical current levels ($\leq 0.5~mA$).
@@
The stringent requirements of a nine-decade dynamic range and a $10~\mu\mathrm{s}$ response time would imply impractically high instantaneous resolution for a direct conversion approach. However, this requirement can be relaxed when considering the limited signal bandwidth of $12~\mathrm{kHz}$ and the variable acquisition time, which can extend up to tens or even hundreds of seconds for non-critical currents. This long-acquisition, high-resolution regime corresponds to beam alignment tracking, where the signals can be considered essentially static. Consequently, the first-order current-mode delta-sigma ADC emerges as an ideal candidate for beam loss monitoring applications. While higher-order modulators offer superior noise shaping and can achieve higher resolution for a given oversampling ratio, their performance advantage diminishes for near-DC signals, where the effective resolution is ultimately limited by the DC gain of the loop filter—a parameter that is comparable to that of a first-order modulator for a given amplifier gain. Moreover, their stability is more complex and less robust, particularly under saturation or overload conditions, where they are prone to prolonged recovery transients or latch-up. In contrast, the first-order architecture provides predictable, unconditional stability and graceful saturation recovery, making it ideally suited for a machine protection system where reliability and deterministic behavior are paramount.
The extreme radiation requirement, exceeding 100 Mrad TID, dictates both technology choice (130 nm CMOS) and design methodology. Digital circuits must be triplicated with majority voting to mitigate single-event upsets (SEUs), while analog blocks require manual enclosed layout and larger guard rings to remain functional after cumulative dose degradation.
@@
Finally, the system’s role as a machine protection element imposes absolute reliability requirements. Redundant communication channels, SEU mitigation, and graceful out-of-range behavior are not optional but mandatory to ensure the BLM system can fulfill its critical function throughout the HL-LHC operational lifetime.
@@ III Converter architecture
The front-end of the modulator consists of an active integrator, which serves two critical purposes: it collects the ionization charge generated by the sensor over time and acts as a filter for the feedback loop. The integrator’s output is continuously monitored by a comparator, clocked at a default frequency of 20 MHz derived by dividing the 80 MHz clock used for the digital circuitry. This frequency represents the optimal trade-off between the integrator’s bandwidth and the analog-to-digital conversion range achievable within the 10 µs response time. On each clock edge, the integrator output is compared to a reference threshold, effecting both sampling and 1-bit quantization. The comparison result drives a time-varying negative feedback loop that prevents the integrator from saturating. This feedback is implemented via a 1-bit digital-to-analog converter (DAC), consisting of a current source and a switch controlled by the comparator’s output. The precision of the compensation charge depends on the stability of the reference current and the clock period. The effectiveness of the sigma-delta conversion relies on the continuous accumulation of the net difference between the input and feedback signals, and on its attenuation by the loop filter gain. A high DC loop gain guarantees that the time-averaged compensation signal closely follows the average input current. To first order, therefore, the achievable resolution of the converter is proportional to the DC gain of the integrator.
The bitstream produced by the quantizer is the digital representation $D_{\textit{out}}$ of the compensation signal whose average value is directly encoded in the density of ’1’s within this bitstream. For static inputs, therefore, the upper bound of the achievable resolution is given by the number of distinguishable codes, expressed as:
@@
where $f_{samp}$ is the sampling frequency and $T_{acq}$ is the acquisition time per measurement. The relationship is depicted in Figure 2
@@
Eq. 1 highlights the intrinsic trade-off between signal bandwidth and conversion resolution characteristic of the sigma-delta topology, making it ideally suited for accurate, high-resolution measurements of slowly varying signals. The bandwidth limitation comes from using longer bitstreams to evaluate the instantaneous average.
@@ III-A Decimation filtering
where M is the bitstream length and n the number of cascaded first-order stages. The case n=1 correspond to a simple moving average, representing the sweet spot for low latency and hardware efficiency, while n=3 provides optimal accuracy. This low-pass filtering also enables the symbol rate to be reduced accordingly without any loss of information.
The inherent trade-off between accuracy and bandwidth offered by sigma-delta conversion can be further exploited by making the outputs of these multiple filter stages available simultaneously. Thereby allowing the BLM system to concurrently use measurement with varying bandwidth and resolution for optimal beam alignment and collimation.
@@ III-B Non-idealities
Beyond the usual limitations of analog components such as noise, offset, slew rate, finite gain-bandwidth, quantizer delay, and switch non-idealities, the converter architecture itself introduces notable non-idealities.
One such phenomenon is the appearance of limit cycles when the static input current approaches a rational multiple of the compensation signal. Under these conditions, the bitstream exhibits periodic regularity, concentrating quantization error energy at specific frequencies in the output spectrum. However, this effect is not critical: as zero-mean tones, their primary impact, if not properly mitigated, is a potential 1-bit oscillation between consecutive output samples.
A second issue arises from transient upsets—such as radiation-induced single-event upsets (SEUs) or out-of-range inputs—that drive the integrator into saturation. During saturation, the bitstream locks to a constant sequence of ’1’s, causing a temporary spill of quantization error into lower frequencies until recovery is complete and normal operation resumes.
Finally although converter overload triggers an emergency beam dump it should be noted that unless an ESD protection discharge occurs, the ionization chamber charge is conserved and eventually absorbed and correctly measured by the converter.
@@
A first-order delta-sigma modulator presents an interesting perspective on stability. From a linear control systems standpoint, the modulator is unconditionally stable: its single-pole loop filter yields a well-behaved step response with no tendency toward oscillation or unbounded growth.
On the other hand, the system can also be viewed as inherently unstable in the sense that a DC input does not produce a DC output. Instead, it generates a continuous high-frequency modulation that, crucially, remains non-periodic for most input values. This non-periodic behavior arises from the sharp nonlinearity introduced by the 1-bit quantizer, whose transfer characteristic is effectively discontinuous (resembling a staircase). Far from being pathological, this irregular oscillation is the fundamental mechanism underlying noise shaping which spreads the quantization error across the upper portion of the spectrum, preventing its concentration into discrete tones.
This apparent paradox stems from the fact that the modulator is not a strictly linear time-invariant (LTI) system as it comprises a nonlinear quantizer and time-varying elements in the switched feedback current. Nevertheless, the system can be approximated as quasi-LTI to a useful extent, allowing feedback theory considerations to remain applicable as a first-order approximation.
@@ IV-A Amplifier
The differential output is realized with a push-pull configuration, biased through a translinear loop to precisely control its quiescent current.
@@ IV-B Quantizer
The 1-bit quantizer, Figure 6, is implemented as a synchronous comparator using a strong-arm latch topology, which is reset during the low phase of the clock and released on its rising edge. The total current drawn from the supply—before the comparator is fully steered—is limited by a tunable bias current programmable from 160 to $640~\mu A$ with 3-bit resolution. This bias current also influences the comparison speed and noise performance.
@@ IV-D DAC
The feedback signal is generated by a 1-bit DAC, which can be trimmed using a 5-bit control word to adjust the converter’s full-scale range according to the expected input current, spanning from $20~\mu A$ to 1.06mA1.06~mA. In order to prevent the modulator from saturating, the feedback current must be at least one LSB larger than the maximum input current.
@@
$\log_{2}\left(\frac{I_{\mathrm{max}}}{I_{\mathrm{noise,rms}}}\right)\approx 10.8~bits$ (3)
This contribution, however, does not limit conversion accuracy in the mid and lower signal ranges. Because the modulator scales a fixed compensation current with the input, the associated noise scales proportionally. As a result, the DAC’s effective noise diminishes with decreasing input current and is further reduced by bitstream averaging. The DAC is implemented as a current mirror using NMOS devices, incorporating cascode transistors and degeneration resistors to enhance noise performance. To prevent radiation-induced mismatch, unused devices are kept active by continuously drawing current from the common-mode voltage (Vcm), rather than being switched off.
@@ V Measurements Results
Post-irradiation performance characterization was carried out on a separate test bench, with annealing occurring only passively during storage at room temperature for a few hours: a scenario representative of realistic operating conditions. No measurable performance degradation was observed; any differences between pre- [13, 12] and post-irradiation measurements were indistinguishable from statistical variation.
Input currents down to 1 pA were generated using a calibrated source-measure unit with guarded triaxial connections (Keithley 6430 equipped with remote amplifier module). Special care was taken in the PCB design to further reduce leakage [13], including the use of shielding and minimizing the distance between connectors and the ASIC to about 2 cm.
@@
Data acquisition was carried out using a BLM crate housing a VFC acquisition board, capable of storing up to 2 Gbit raw bit-stream per measurement. The acquired data were then exported to a PC for post-processing. Each measurement corresponds to a fixed DC input current. The equivalent noise current is evaluated as a function of acquisition time by subdividing the raw bitstream into smaller, consecutive segments. For each segment, the average density of ’1’s is computed, and the noise is then taken as the standard deviation of these repeated measurements. The equivalent input referred noise current assumes Figure 11 reports the equivalent input-referred noise as a function of acquisition time for input currents spanning from 1 mA down to 10 pA. The most relevant measurement combinations are detailed in Table II indicates that sub-pA accuracy requires a minimum of ten seconds of integration and input currents below 10 nA. The dominant noise source shifts with current levels; as predicted by simulations, the compensation DAC limits accuracy at higher current ranges, while the operational amplifier dominates at lower ranges. Grayed-out cells in the table indicate incompatible parameter combinations for which the noise exceeds the targeted measurement range. By relating the maximum static input current to the minimum noise level observed over a 100 s integration window, the resulting operating dynamic range is:
$DR_{oper}=20\cdot log_{10}\left(\frac{I_{in,~max}}{i_{noise,~min}}\right)\approx 201.94~dB.$ (4)
@@
However, this value is only attainable across multiple input ranges, as the noise contribution from the compensation current scales with the square root of the density of ’1’s in the bitstream. Table III presents the measured ENOB as a function of the integration time and input current. With a $10~\mu s$ integration window, 11-bit effective resolution is achieved only for the highest current. At 100 ms, however, 15-bits accuracy extends over three decades, from $10~\mu A$ to 1 mA. For 100 s, 15 bits are exceeded for currents spanning from 1 mA down to 10 nA.
@@
The ADC transfer characteristic and its integral non-linearity (INL) were evaluated for the fastest acquisition setting (Figure 12). With a 20 MHz sampling frequency, this configuration yields 200 distinct output codes within the $10~\mu s$ integration window. Over the corresponding input range, $\left[1~mA,~5~\mu A\right]$, the uncalibrated INL remains within +0.2% and -0.3% of full scale in compliance with the BLM application requirements.
@@
Figure 11: Equivalent input-referred noise current versus acquisition time. Each point represents the standard deviation of multiple repeated measurements, with the acquisition time determined by the length of the averaged bitstream.
@@
TABLE II: RMS noise current vs input currents and integration times
Input Current Integration Time
@@
100 nnA 200 nnA 60 nnA nnA 600 ppA 60 ppA ppA ppA 1.8 ppA
10 nnA 60 nnA 20 nnA nnA 600 ppA 60 ppA ppA 0.6 ppA 130 ffA
nnA 20 nnA nnA nnA 600 ppA 60 ppA ppA 0.6 ppA 130 ffA
100 ppA 10 nnA nnA nnA 300 ppA 60 ppA ppA 0.6 ppA 80 ffA
10 ppA nnA nnA 600 ppA 200 ppA 60 ppA ppA 0.6 ppA 80 ffA
@@
Input Current $ENOB(T_{int})$

### v2

# A 200 dB Dynamic Range Radiation‑Hard Delta‑Sigma Current Digitizer for Beam Loss Monitoring

@@ Abstract
This manuscript describes a radiation-hardened current-mode delta-sigma ADC fabricated in a standard 130 nm CMOS technology and qualified for total ionizing doses up to 100 Mrad. The operational signal range achieved with 100 s integration window exceeds 200 dB. The converter is designed for beam loss monitoring applications in high-energy physics, where it must handle input currents spanning nine decades, from 1 mA down to 1 pA, while providing a fast 10 µs response time for machine protection. To meet these conflicting requirements, the architecture exploits the inherent trade-off between resolution and acquisition time provided by delta-sigma conversion: a first-order architecture, sampling at 20 MHz, delivers 11 bit effective resolution within the critical 10 µs window for critical currents around 1 mA. Integration times above 10 s enable the sub-picoampere resolution required for precise beam alignment and background monitoring. The chip integrates two independent channels, consumes 25 mW from a 1.2 V supply, and relies on radiation-hardening techniques such as triple-redundant digital logic, custom ESD protections and manual enclosed layout for critical analog transistors. Post-irradiation measurements up to 100 Mrad show no significant performance degradation, and the uncalibrated integral nonlinearity remains within [+4,–5] LSBs over the 1 mA to 5 µA range. The converter’s flexibility and radiation tolerance make it suitable not only for the HL-LHC beam loss monitoring upgrade but also for other precision current measurement applications in harsh environments.
@@ I Introduction
RADIATION-TOLERANT current digitizers primarily find their use in harsh environments where radiation monitoring is required, like in high energy physics experiments and nuclear power plants. For example, at CERN the radiation levels are carefully monitored around the Large Hadron Collider (LHC) in order to protect the machine components and to accurately position the absorber jaws of the mechanical beam collimation system. In both cases the sensing element is a gas filled ionization chamber [1, 2, 3] that produces a fixed charge rate when subject to constant radiation levels; the corresponding electrical signal ranges from 1 mA to few picoamperes (pA) which represent respectively the maximum tolerable radiation energy, absorbed within the response time of the beam dumping system, and the typical background beam losses. Various current sensing interfaces exist either built with discrete components [4] or as integrated circuits [5] but their radiation tolerance limits impose a lower bound for the length of the cable connecting the interface with the sensor. Long cables limit the detection dynamic range because of to interference, noise charge associated with the parasitic capacitance and leakage currents. The radiation levels around the LHC are maximum at the collimators preceding the interaction points (IP3 and IP7) [6] and are expected to grow after the High Luminosity upgrade foreseen by 2025. The radiation around the collider originates from the losses of the beams circulating inside the machine, since the relevant quantity is the dose rate absorbed within one turn rather than its instantaneous value, the application band of interest spans from dc up to 12 kHz. An integrated radiation tolerant frontend has therefore been developed to comply with the stringent requirements of radiation monitoring in harsh environments. To this end a standard CMOS 130 nm technology was selected after being successfully characterized to tolerate a total ionizing dose (TID) of 200 Mrad [7, 8, 9] when adopting specific radiation hardness measures to prevent latch up, drain/source leakage and to tolerate single event upsets (SEU) and transients (SET). The remainder of this manuscript is organized as follows. Section II presents the system requirements and design drivers derived from the Beam Loss Monitoring (BLM) application and the HL-LHC upgrade constraints. Section III describes the architecture of the proposed delta-sigma current digitizer, discussing its resolution limit and stability considerations while also providing the foundational concepts for readers less familiar with this architecture. Section IV details the circuit-level implementation, including the integrator, quantizer, adaptive clock control and DAC. Section V reports the experimental results, including noise performance, linearity, and dynamic range as a function of acquisition time, for samples irradiated with X-rays up to 100 Mrad. Finally, Section VI concludes the paper.
@@ II System Requirements and Design Drivers
The BLM system [10] is a critical protection element of the LHC accelerator complex, designed to prevent quenches of superconducting magnets and equipment damage caused by excessive radiation following beam misalignment. The system continuously measures radiation levels along the accelerator using approximately 4000 ionization chambers installed throughout the 27 km of the circular collider. When ionizing radiation passes through the gas-filled chambers, the resulting charge is extracted thanks to a high-voltage bias (1.5 kV) and converted into a current proportional to the instantaneous beam loss. In the original LHC design these signals are transmitted via coaxial cables, up to 800 m in length, to about 750 acquisition modules for processing and analysis.
To ensure compatibility with the higher luminosity and correspondingly increased radiation levels expected from the HL-LHC upgrade, a new generation of front-end electronics must meet stringent performance, reliability, and environmental constraints. The enhanced radiation tolerance enables the readout electronics to be installed closer to the sensors, thereby improving signal integrity. However, depending on the radiation levels and associated aging effects, the electronics must also remain compatible with cable lengths of up to 50 m to preserve installation flexibility, although with relaxed requirements for the non-critical current levels ($\leq 0.5$ mA).
@@
The stringent requirements of a nine-decade dynamic range and a $10~\mu\mathrm{s}$ response time would imply impractically high instantaneous resolution for a direct conversion approach. However, this requirement can be relaxed when considering the limited signal bandwidth of $12~\mathrm{kHz}$ and the variable acquisition time, which can extend up to tens or even hundreds of seconds for less critical currents. This long-acquisition, high-resolution regime corresponds to beam alignment tracking, where the signals can be considered essentially static. Consequently, the first-order current-mode delta-sigma ADC emerges as an ideal candidate for beam loss monitoring applications. While higher-order modulators offer superior noise shaping and can achieve better resolution for a given oversampling ratio, their performance advantage diminishes at lower frequencies, where the effective resolution is ultimately limited by the DC gain of the loop filter, namely a parameter that is independent on the converter’s order. Moreover, their stability is harder to achieve and less robust to perturbations, particularly under saturation or overload conditions, where they are prone to prolonged recovery transients or even deadlock. In contrast, the first-order architecture provides predictable, unconditional stability and graceful saturation recovery, making it ideally suited for a machine protection system where reliability and deterministic behavior are paramount.
The extreme radiation requirement, exceeding 100 Mrad total integrated dose (TID), dictates both technology choice (130 nm CMOS) and design methodology. Digital circuits must be triplicated with majority voting to mitigate SEUs, while analog blocks require custom enclosed layout and more stringent latch-up constraints to preserve functionality and performance after cumulative dose degradation.
@@
Finally, the system’s role as a machine protection element imposes absolute reliability requirements. Redundant communication channels, SET resilience, and graceful out-of-range behavior are not optional but mandatory to ensure the BLM system can fulfill its critical function throughout the HL-LHC operational lifetime.
@@ III Converter architecture
[…]
The bitstream produced by the quantizer is the digital representation $D_{\textit{out}}$ of the compensation signal whose average value is directly encoded in the density of ’1’s within this bitstream. For static inputs, therefore, the upper bound under ideal 1 bit averaging of the achievable resolution corresponds to the number of distinguishable codes, expressed as:
@@
where $f_{samp}$ is the sampling frequency and $T_{acq}$ is the acquisition time per measurement.
@@
Eq. 1 highlights the intrinsic trade-off between signal bandwidth and conversion resolution characteristic of the sigma-delta topology, making it ideally suited for accurate, high-resolution measurements of slowly varying signals. The bandwidth limitation comes from using longer bitstreams to evaluate the instantaneous average. The relationship is depicted in Figure 2
@@ III-A Decimation filtering
where M is the bitstream length and n the number of cascaded first-order filters. The case n=1 corresponds to a simple moving average, representing the sweet spot for low latency and hardware efficiency, while n=3 provides optimal accuracy. This low-pass filtering also enables the symbol rate to be reduced accordingly without any loss of information.
The inherent trade-off between accuracy and bandwidth offered by sigma-delta conversion can be further exploited by making the outputs of these multiple filter stages available simultaneously. Thereby allowing the BLM system to concurrently use measurements with varying bandwidth and resolution for optimal beam alignment and collimation.
@@ III-B Non-idealities
Beyond the usual limitations of analog components such as noise, offset, slew rate, finite gain-bandwidth, quantizer delay, and switch non-idealities, the converter architecture itself introduces notable deficiencies.
One such phenomenon is the appearance of limit cycles due to the static input signal being a rational multiple of the compensation current. Under these conditions, the bitstream exhibits periodic regularity, concentrating quantization error energy at specific frequencies in the output spectrum. However, this effect is not critical: as zero-mean tones, their primary impact, if not properly mitigated, is a potential 1-LSB oscillation between consecutive output samples.
A second issue arises from transitory perturbations, such as SEUs or out-of-range inputs, that drive the integrator into saturation. During saturation, the bitstream locks to a constant sequence of ’1’s, causing a temporary spill of quantization error into lower frequencies until recovery is complete and normal operation resumes. However, this converter overload does not prevent the detection of critical radiation levels and the initiation of an emergency beam dump. Moreover, unless an ESD protection discharge occurs, the charge produced by the ionization chamber is conserved and its average rate is correctly measured by the converter.
@@
A first-order delta-sigma modulator offers an unconventional perspective on stability. The loop contains an ideal discrete-time integrator with a pole at $z=1$, which is marginally stable: in the absence of feedback, a step input produces an output that grows unbounded, albeit linearly rather than exponentially. The boundedness of the modulator is provided by the action of the nonlinear feedback implemented through the 1 bit quantizer and DAC.
This nonlinear feedback forces polarity reversals whenever the integrator state crosses the decision threshold, preventing monotonic growth of the state variable. For input signals within the available feedback range, the system does not converge to a static equilibrium but instead evolves toward a bounded oscillatory trajectory in state-space. In this sense, the modulator behaves similarly to a relaxation oscillator whose amplitude is constrained by feedback.
From a purely linear viewpoint, the sustained oscillatory behavior may resemble that of a feedback system with insufficient phase margin. However, the underlying mechanism is fundamentally different: the oscillation is not caused by linear pole placement but by the discontinuous switching action of the quantizer. It is therefore intrinsic to the intended operation of the converter rather than a symptom of linear instability.
The staircase-like nonlinearity of the quantizer and DAC introduces irregular switching patterns for most input values. This irregularity is the mechanism that spreads quantization error energy toward higher frequencies, enabling noise shaping and preventing its concentration into discrete tones, provided pathological limit cycles are avoided.
Despite its nonlinear nature, the modulator can be approximated as quasi–linear time-invariant for analytical purposes. By linearizing the quantizer and DAC around their average operating point, classical feedback theory can be applied as a first-order approximation, yielding useful insight into loop gain, signal and noise transfer functions, and stability margins within the intended operating range. When the input exceeds the available feedback current, the state escapes this bounded trajectory and grows until physical saturation occurs; normal bounded operation resumes once the input returns within its operational range.
@@ IV-A Amplifier
The differential output is realized with a push-pull configuration, biased through a translinear loop to precisely control its quiescent current. The class-AB configuration prevents slew-rate limitations during large-signal excursions without incurring excessive power penalty.
@@ IV-B Quantizer
The 1 bit quantizer, Figure 6, is implemented as a synchronous comparator using a strong-arm latch topology, which is reset during the low phase of the clock and released on its rising edge. The total current drawn from the supply—before the comparator is fully steered—is limited by a tunable bias current programmable from 160 to $640~\mu A$ with 3 bit resolution. This bias current also influences the comparison speed and noise performance.
@@ IV-D DAC
The feedback signal is generated by a 1 bit DAC, which can be trimmed using a 5 bit control word to adjust the converter’s full-scale range according to the expected input current, spanning from $20~\mu A$ to 1.06mA1.06~mA. In order to prevent the modulator from saturating, the feedback current must be at least one LSB larger than the maximum input current.
@@
$Resolution_{DAC}=\log_{2}\left(\frac{I_{\mathrm{max}}}{I_{\mathrm{noise,rms}}}\right)\approx 10.8~bits$ (3)
This contribution limits only the conversion of full-scale signals, as it is not injected at every sampling period for lower currents. Consequently, it does not impair accuracy in the mid and lower signal ranges. Since the modulator scales a fixed compensation current with the input, the associated noise scales proportionally. As a result, the DAC’s effective noise diminishes with decreasing input current and is further reduced by bitstream averaging.
The DAC is implemented as a current mirror using NMOS devices, incorporating cascode transistors and degeneration resistors to enhance noise performance. To prevent radiation-induced mismatch, unused devices are kept active by continuously drawing current from the common-mode voltage (Vcm), rather than being switched off.
@@ V Measurements Results
Post-irradiation performance characterization was carried out on a separate test bench, with annealing occurring only passively during storage at room temperature for a few hours: a scenario representative of realistic operating conditions. No measurable performance degradation was observed; any differences between pre- [13, 12] and post-irradiation measurements were indistinguishable from statistical variation. Single-event upset (SEU) resilience tests under heavy-ion and neutron irradiation have not yet been performed due to resource constraints. However, no significant deviations are expected, as the design strategies employed are consistent with those used in other functionally similar ASICs that have been successfully qualified and integrated into LHC’s detectors. Input currents down to 1 pA were generated using a calibrated source-measure unit with guarded triaxial connections (Keithley 6430 equipped with remote amplifier module). Special care was taken in the PCB design to further reduce leakage [13], including the use of shielding and minimizing the distance between connectors and the ASIC to about 2 cm.
@@
Data acquisition was carried out using a BLM crate housing a VFC acquisition board, capable of storing up to 2 Gbit raw bit-stream per measurement. The acquired data were then exported to a PC for post-processing. Each measurement corresponds to a fixed DC input current. The equivalent noise current is evaluated as a function of acquisition time by segmenting the raw output into consecutive, non-overlapping bitstream intervals. For each segment, the average density of ’1’s is computed using a rectangular time window, which corresponds to a sinc-shaped frequency response. The noise is then taken as the standard deviation of these repeated measurements. The resulting input-referred noise current assumes a perfectly linear (or calibrated) transfer characteristic.
Figure 11 reports the equivalent input-referred noise as a function of acquisition time for input currents spanning from 1 mA down to 10 pA. The most relevant measurement combinations are detailed in Table II: sub-pA accuracy can be reached under, at least, ten seconds of integration and input currents below 10 nA. The dominant noise source shifts with input levels; as predicted by simulations, the compensation DAC limits accuracy at higher current ranges, while the operational amplifier dominates at lower ranges. Grayed-out cells in the table indicate incompatible parameter combinations for which the noise exceeds the targeted measurement range. The last row reports the operating dynamic range corresponding to each integration window and defined as:
$DR_{oper}=20\cdot log_{10}\left(\frac{I_{in,~max}}{i_{noise,~min}}\right)$ (4)
For a 100 s integration window, this dynamic range reaches 201.9 dB. Conversely, Table III reports the effective number of bits (ENOB) for each current, measured with the minimum, mid and maximum integration windows, according to:
$ENOB(I_{in},T_{acq})=log_{2}\left(\frac{I_{in}}{i_{noise}(T_{acq})}\right).$ (5)
@@
With a $10~\mu s$ integration window, 11 bit effective resolution is achieved only for the highest current. At 100 ms, however, 15 bit accuracy extends over three decades, from 1 mA to $10~\mu A$. For 100 s, 15 bit ENOB is exceeded for currents spanning from 1 mA down to 10 nA.
@@
Figure 12 presents the measured ADC transfer characteristic and integral non-linearity (INL) for the fastest acquisition setting of $10~\mu\mathrm{s}$; sampling at 20 MHz can only produce 200 distinct output codes. Across the corresponding input range of $1~\mathrm{mA}$ to $5~\mu\mathrm{A}$, the uncalibrated INL stays within $0.2\%$ and $0.3\%$ of full scale, complying with BLM specifications.
@@
Figure 11: Equivalent input-referred noise current as a function of integration time. The noise is obtained from the standard deviation of consecutive measurements, each corresponding to the averaged bitstream over the indicated duration.
@@
TABLE II: Measured RMS Noise Current vs. Input Current $I_{\text{in}}$ and Integration Time $T_{\text{acq}}$
@@
100 nA 174 nA 51 nA 4.6 nA 598 pA 31 pA 5.2 pA 2.1 pA 1.6 pA
10 nA 57 nA 18 nA 5.1 nA 552 pA 51 pA 4.4 pA 515 fA 134 fA
1 nA 21 nA 6.8 nA 2.1 nA 555 pA 46 pA 3.9 pA 530 fA 123 fA
100 pA 9.2 nA 2.9 nA 922 pA 283 pA 56 pA 5.7 pA 575 fA 81 fA
10 pA 6.1 nA 2.0 nA 613 pA 191 pA 52 pA 4.9 pA 566 fA 77 fA
[…]
   dB 66.1 85.9 105.9 125.1 145.0 166.2 184.9 202.3
   bits 11.0 14.3 17.6 20.8 24.1 27.6 30.7 33.6
Gray cells indicate configurations where the measured noise exceeds the nominal input current and are therefore not included in the $DR_{oper}$ estimation.
@@
Input Current $ENOB(T_{acq})$